# Neural Enhancement of the Traditional Wang–Sheeley–Arge Solar Wind Relation

### v1

@@ Abstract
The Wang–Sheeley–Arge (WSA) model has been the cornerstone of operational solar wind forecasting for nearly two decades, owing to its simplicity and physics-based formalism. However, its performance is strongly dependent on several empirical parameters that are typically fixed or tuned manually, limiting its adaptability across varying solar conditions. In this study, we present a neural enhancement to the WSA framework (referred to as WSA+) that systematically optimizes the empirical parameters of the WSA solar wind speed relation using in-situ observations within a differentiable physics-constrained pipeline. The approach operates in two stages: first, a neural optimizer adjusts WSA parameters independently for each Carrington Rotation to better match the observed solar wind data. Then, a neural network learns to predict these optimized speed maps directly from magnetogram-derived features. This enables generalization of the optimization process and allows inference for new solar conditions without manual tuning. The developed WSA+ preserves the interpretability of the original relation while significantly improving the match with OMNI in-situ data across multiple performance metrics, including correlation and error statistics. It consistently outperforms the traditional WSA relation across both low and high solar activity periods, with average improvements of approximately 40 percent. By integrating data-driven learning with physical constraints, WSA+ offers a robust and adaptable enhancement, with immediate utility as a drop-in replacement in global heliospheric modeling pipelines.
@@ 1 Introduction
Space weather events driven by solar activity can severely impact satellites, aviation, communication systems, and electrical infrastructure on Earth. Among the various drivers, the solar wind is a continuous structured outflow of plasma from the Sun. It serves as a persistent background condition that modulates heliospheric dynamics and preconditions the near-Earth environment. It also strongly influences coronal mass ejection (CME) propagation (P. Mayank et al., 2023; C. Kay et al., 2024) and CME-CME interactions (P. Mayank et al., 2024; S. V. Thampi et al., 2025). Accurate modeling of the ambient solar wind is therefore essential for both research and operational space weather forecasting.
@@
In this work, we introduce a hybrid machine learning (ML) framework that overcomes this constraint by first optimizing WSA parameters per CR using a neural optimizer, and then training a neural network to generalize across CRs. The surrogate model predicts full 2D solar wind speed maps at 21.5 solar radius (0.1 AU), capturing structural variability while preserving physical interpretability. This hybrid ML approach allows to address the limitation of tunable WSA parameters through learning the generalized pattern of their optimized set of values across the solar activity.
The structure of this paper is as follows: in Section 2, we describe the two-stage modeling pipeline, including the neural optimizer and the neural network. In Section 3, we present performance results across 129 Carrington Rotations and analyze spatial and time series validation. Section 4 discusses implications for operational forecasting and limitations of the current approach. We conclude in Section 5 with a summary and outlook.
@@ 2 Methodology
The Wang–Sheeley–Arge (WSA) model computes solar wind speed by relating it to magnetic topology derived from observed synoptic magnetograms. This mapping relies on the Potential Field Source Surface (PFSS) model, which assumes a current-free inner-corona and solves the elliptic Laplace equation to compute the scalar potential of the magnetic field (M. D. Altschuler & G. Newkirk, 1969). From this solution, the flux-tube expansion factor at the source surface ($f_{s}$) and the minimum angular distance ($D$) to the coronal hole boundary are derived, both of which serve as key inputs to the empirical WSA wind speed relation:
@@
Figure 1: Schematic diagram of the undertaken method to optimize and generalize the WSA solar wind speed maps.
@@ 2.1 Neural Optimizer
To address the limitations of fixed empirical parameters, we develop a differentiable optimization framework that couples the WSA speed relation with the Heliospheric Upwind Extrapolation (HUX) model to produce wind speed at 1 AU (P. Riley & R. Lionello, 2011). The empirical parameters in the WSA relation are treated as free variables constrained within physically plausible bounds around their default values (see Appendix A). These parameters are optimized on a per–CR basis by minimizing a composite loss. This loss is computed from a set of statistical metrics that quantify the mismatch between predicted and observed solar wind speeds at L1, including root mean square error (RMSE), mean absolute error (MAE), dynamic time warping (DTW), and Pearson correlation coefficient (PCC).
This loss is minimized using the Limited-memory Broyden–Fletcher–Goldfarb–Shanno (L–BFGS) algorithm (D. C. Liu & J. Nocedal, 1989), a quasi-Newton method that approximates the inverse Hessian matrix to efficiently navigate the parameter space. Its low memory footprint and fast convergence make it particularly well-suited for optimizing low-dimensional problems with smooth, differentiable objective functions—like our case, where the number of empirical WSA parameters is small and the loss surface is well-behaved. The resulting optimized parameter maps capture CR-specific tuning, enabling improved accuracy without sacrificing interpretability.
@@ 2.2 Neural Network
While per-CR optimization significantly improves the alignment between predicted and observed solar wind speeds, it requires in-situ data and is therefore not directly usable for forecasting. To generalize this optimization step and enable predictive capability, we train a neural network to infer the optimized WSA parameters directly from magnetogram based PFSS–derived features. Figure 1 demonstrates all the critical steps for the development of WSA+ model.
The input to the model consists of 2D maps of the expansion factor and minimum angular distance, while the targets are the WSA optimized maps obtained from the neural optimization stage. We adopt a Swin Transformer (Z. Liu et al., 2021) architecture for this task, leveraging its ability to capture multi-scale spatial correlations and geometric context from the input maps. By learning from the broader distribution of optimized maps, the network produces more realistic and smoothly varying parameter predictions, improving both the fidelity and generalizability of the solar wind speed maps derived from the WSA formulation. More information on the implemented Swin transformer is in Appendix B.
@@ 3 Results
The performance of WSA+ is assessed through both spatial and temporal comparisons against the default WSA configuration and its CR-specific optimized variant. Following subsections demonstrate the results for the global solar wind speed structure through 2D maps, followed by in-situ time-series validation at the L1 point.
@@ 3.2 In-situ Analysis
Consistently across CRs, WSA+ demonstrates intermediate performance between the default and CR-specific optimized profiles. Figure 4 summarizes the comparison metrics between the default WSA and WSA+ outputs against in-situ observations at L1, across considered CRs. WSA+ shows substantial improvements over the default WSA: a 60.5% reduction in DTW, 45.9% in MAE, and 40.6% in RMSE, along with a marginal increase of 1.3% in PCC. This pattern of significant gains in error-based metrics, with minimal change in linear correlation, suggests that WSA+ effectively learns the underlying empirical relationship with $f_{s}$ and $D$, while preserving the structural integrity of the WSA formulation.
@@ 5 Conclusion
The resulting model, WSA+, shows an overall improvement of more than 40% over the traditional WSA model. It serves as a direct, physically consistent enhancement to the WSA pipeline, and is made openly available through a Python package: wsaplus (https://pypi.org/project/wsaplus/). Designed for operational usability, WSA+ can be seamlessly integrated into existing space weather modeling frameworks as a drop-in replacement for default WSA outputs, without requiring additional inputs or tuning.
@@
This research was supported by the NASA Living with a Star Jack Eddy Postdoctoral Fellowship Program, administered by UCAR’s Cooperative Programs for the Advancement of Earth System Science (CPAESS) under award #80NSSC22M0097. This work was partially supported by NASA under awards No 80NSSC23M0192, 80NSSC20K1580, 80NSSC21K1555. Data Availablity Statement: The used GONG synoptic magnetograms maps can be freely obtained from https://gong.nso.edu/data/magmap/crmap.html. The OMNI data are taken from the Goddard Space Flight Center, accessible at https://spdf.gsfc.nasa.gov/pub/data/omni/. The used CME catalog can be accessed from https://wind.nasa.gov/ICME_catalog/ICME_catalog_viewer.php. The dataset presented in this paper, along with the final model checkpoint, are archived at https://doi.org/10.5281/zenodo.16883042. Code Sharing Statement: All the developed codes, including training, testing and visualization, can be accessed through https://github.com/PrateekMayank/wsaplus. This appendix provides technical details regarding the parameter optimization method and neural network architecture used in this study. Section A outlines the empirical parameter ranges and fitting strategy. Section B describes the WSA+ model architecture based on Swin Transformer, along with training diagnostics.
@@ Appendix B Model Architecture and Training
Figure 7 compares the model performance across training, validation, and test sets using five evaluation metrics. WSA+ consistently outperforms the default WSA–HUX pipeline in RMSE, MAE, and DTW, with improvements sustained across all dataset splits. While the improvement in PCC is not consistent, the overall FIT score is higher for WSA+ (0.71 test score versus 0.49 from default WSA), demonstrating reliable predictive performance. A summary of these statistics is also provided within Figure 7.

### v2

@@ Abstract
The Wang–Sheeley–Arge (WSA) model has been the cornerstone of operational solar wind forecasting for nearly two decades, owing to its simplicity and physics-based formalism. However, its performance is strongly dependent on several empirical parameters that are typically fixed or tuned manually, limiting its adaptability across varying solar conditions. In this study, we present a neural enhancement to the WSA framework (referred to as WSA+) that systematically optimizes the empirical parameters of the WSA solar wind speed relation using in-situ observations within a differentiable physics-constrained pipeline. The approach operates in two stages: first, a neural optimizer adjusts WSA parameters independently for each Carrington Rotation to better match the observed solar wind data. Then, a neural network learns to predict these optimized speed maps directly from magnetogram-derived features. This enables generalization of the optimization process and allows inference for new solar conditions without manual tuning. WSA+ preserves the interpretability of the original relation while significantly improving the match with OMNI in-situ data across multiple performance metrics, including correlation and error statistics. It consistently outperforms the traditional WSA relation across both low and high solar activity periods, with average improvements of approximately 40 percent. By integrating data-driven learning with physical constraints, WSA+ offers a robust and adaptable enhancement, with immediate utility as a drop-in replacement in global heliospheric modeling pipelines.
[…]
@@ 1 Introduction
Space weather events driven by solar activity can severely impact satellites, aviation, communication systems, and electrical infrastructure on Earth. Among the various drivers, the solar wind is a continuous structured outflow of plasma from the Sun. It serves as a persistent background condition that modulates heliospheric dynamics and preconditions the near-Earth environment. It also strongly influences coronal mass ejection (CME) propagation (R. M. Winslow et al., 2021; P. Mayank et al., 2023; C. Kay et al., 2024) and CME-CME interactions (N. Lugaz et al., 2017; P. Mayank et al., 2024; S. V. Thampi et al., 2025). Accurate modeling of the ambient solar wind is therefore essential for both research and operational space weather forecasting.
@@
In this work, we introduce a hybrid machine learning (ML) framework that overcomes this constraint by first optimizing WSA parameters per CR using a neural optimizer, and then training a neural network to generalize across CRs. The surrogate model predicts full 2D solar wind speed maps at 21.5 solar radius (0.1 AU), capturing structural variability while preserving physical interpretability. This hybrid ML approach allows us to address the limitation of tunable WSA parameters through learning the generalized pattern of their optimized set of values across the solar cycle.
The structure of this paper is as follows: in Section 2, we describe the two-stage modeling pipeline, including the neural optimizer and the neural network. In Section 3, we present performance results across 129 CRs and analyze spatial and time series validation. Section 4 discusses implications for operational forecasting and limitations of the current approach. We conclude in Section 5 with a summary and outlook.
@@ 2 Methodology
The WSA model computes solar wind speed by relating it to open fieldlines derived from observed synoptic magnetograms. This mapping relies on the Potential Field Source Surface (PFSS) model, which assumes a current-free inner-corona and solves the elliptic Laplace equation to compute the scalar potential of the magnetic field (M. D. Altschuler & G. Newkirk, 1969). From this solution, the flux-tube expansion factor at the source surface ($f_{s}$) and the minimum angular distance ($D$) to the coronal hole boundary are derived, both of which serve as key inputs to the empirical WSA wind speed relation:
@@
Figure 1: Schematic diagram of the undertaken method to optimize and generalize the WSA solar wind speed maps. PFSS is used to derive expansion factor and minimum angular distance maps from synoptic magnetograms. These 2D maps serve as input for both neural optimizer and neural network. L-BFGS scheme is employed to generate optimized 2D WSA maps by fitting the WSA parameters to OMNI data. Swin Transformer based neural network is trained on these optimized WSA maps to learn how to produce them directly from the PFSS-derived input maps.
@@ 2.1 Neural Optimizer
To address the limitations of fixed empirical parameters, we develop a differentiable optimization framework that couples the WSA speed relation with the Heliospheric Upwind Extrapolation (HUX; P. Riley & R. Lionello, 2011) model to produce wind speed at 1 AU. The empirical parameters in the WSA relation are treated as free variables constrained within physically plausible bounds around their default values (see Appendix A). These parameters are optimized on a per–CR basis by minimizing a composite loss. This loss is computed from a set of statistical metrics that quantify the mismatch between predicted and observed solar wind speeds at L1, including root mean square error (RMSE), mean absolute error (MAE), dynamic time warping (DTW), and Pearson correlation coefficient (PCC).
This loss is minimized using the Limited-memory Broyden–Fletcher–Goldfarb–Shanno (L–BFGS) algorithm (D. C. Liu & J. Nocedal, 1989), a quasi-Newton method that approximates the inverse Hessian matrix to efficiently navigate the parameter space. Its low memory footprint and fast convergence make it particularly well-suited for optimizing low-dimensional problems with smooth, differentiable objective functions—like our case, where the number of empirical WSA parameters is small and the loss surface is well-behaved.
The resulting set of per-CR best-fit parameters is used to generate the optimized 2D WSA speed maps. These maps act as the ground truth for the neural network.
@@ 2.2 Neural Network
While per-CR optimization significantly improves the alignment between predicted and observed solar wind speeds, it requires in-situ data and is therefore not directly usable for forecasting. To generalize this optimization step and enable predictive capability, we train a neural network to infer the optimized WSA maps directly from magnetogram based PFSS–derived features. Figure 1 demonstrates all the critical steps for the development of the WSA+ model.
The input to the model consists of 2D maps of the expansion factor ($f_{s}$) and minimum angular distance ($D$), while the targets are the optimized WSA maps obtained from the neural optimization stage. Thus, the model is trained solely on the optimized 2D WSA maps — not on 1D OMNI time series data or the best-fit numerical values of the seven WSA parameters.
We adopt a Swin Transformer (Z. Liu et al., 2021) architecture for this task, leveraging its ability to capture multi-scale spatial correlations and geometric context from the input maps. By learning from the 14 years of broad distribution of optimized maps, the network generates more realistic and smoothly varying output maps, in contrast to the parametrically optimized WSA maps, which are often overfitted to sub-Earth latitudes. This neural approach improves both the fidelity and generalizability of the solar wind speed maps derived from the WSA formulation. More information on the implemented Swin Transformer is in Appendix B.
@@ 3 Results
The performance of WSA+ is assessed through both spatial and temporal comparisons against the default WSA configuration and its CR-specific optimized variant. The following subsections demonstrate the results for the global solar wind speed structure through 2D maps, followed by in-situ time-series validation at the L1 point.
@@ 3.2 In-situ Analysis
Consistently across CRs, WSA+ demonstrates intermediate performance between the default and CR-specific optimized profiles. Figure 4 summarizes the comparison metrics between the default WSA and WSA+ outputs against in-situ observations at L1, across considered CRs. The train, validation, and test datasets are shown in different colors. For the entire dataset, WSA+ shows substantial improvements over the default WSA: a 60.5% reduction in DTW, 45.9% in MAE, and 40.6% in RMSE, along with a marginal increase of 1.3% in PCC. This pattern of significant gains in error-based metrics, with minimal change in linear correlation, suggests that WSA+ effectively learns the underlying empirical relationship with $f_{s}$ and $D$, while preserving the structural integrity of the WSA formulation.
@@ 5 Conclusion
The resulting model, WSA+, shows an overall improvement of 40% across the entire dataset (39% on the held-out test set) over the traditional WSA model. It serves as a direct, physically consistent enhancement to the WSA pipeline, and is made openly available through a Python package: wsaplus (https://pypi.org/project/wsaplus/). Designed for operational usability, WSA+ can be seamlessly integrated into existing space weather modeling frameworks as a drop-in replacement for default WSA outputs, without requiring additional inputs or tuning.
@@
This research was supported by the NASA Living with a Star Jack Eddy Postdoctoral Fellowship Program, administered by UCAR’s Cooperative Programs for the Advancement of Earth System Science (CPAESS) under award #80NSSC22M0097. This work was partially supported by NASA under awards No 80NSSC23M0192, 80NSSC20K1580, 80NSSC21K1555. Data Availablity Statement: The used GONG synoptic magnetograms maps can be freely obtained from https://gong.nso.edu/data/magmap/crmap.html. The OMNI data are taken from the Goddard Space Flight Center, accessible at https://spdf.gsfc.nasa.gov/pub/data/omni/. The used CME catalog can be accessed from https://wind.nasa.gov/ICME_catalog/ICME_catalog_viewer.php. The dataset presented in this paper, along with all the codes and the final model checkpoint, are archived at https://doi.org/10.5281/zenodo.17309172. Code Sharing Statement: All the developed codes, including training, testing and visualization, can be accessed through https://github.com/PrateekMayank/wsaplus. This appendix provides technical details regarding the parameter optimization method and neural network architecture used in this study. Section A outlines the empirical parameter ranges and fitting strategy. Section B describes the WSA+ model architecture based on Swin Transformer, along with training diagnostics.
@@ Appendix A Optimization of Empirical Parameters
This behavior, where best-fit parameter values frequently reach the bounds of their prescribed ranges, has also been observed in prior studies by P. Riley et al. (2015) and O. Issan et al. (2023). P. Riley et al. (2015) optimized five WSA parameters against ACE/WIND observations using the PCC metric, while O. Issan et al. (2023) used Bayesian inference with Markov Chain Monte Carlo to estimate the distribution of five most influential WSA parameters. These findings, consistent with our results, likely reflect the limited identifiability of WSA parameters when tuned using 1D in-situ time series data within physically constrained bounds.
@@ Appendix B Model Architecture and Training
Figure 7 compares the model performance across training, validation, and test sets using five evaluation metrics. WSA+ consistently outperforms the default WSA–HUX pipeline in RMSE, MAE, and DTW, with improvements sustained across all dataset splits. While the improvement in PCC is not consistent, the overall FIT score is consistently higher for WSA+ across all datasets, demonstrating reliable predictive performance. However, the ratio of metric values of WSA and WSA+ is slightly different for training, validation and test sets. This variation arises primarily because the ratio depends on which CRs are included in each dataset. Since the fidelity of the GONG+PFSS+WSA+HUX pipeline varies from CR to CR, the relative performance metrics will also exhibit similar variation. A summary of these statistics is also provided within Figure 7.